The current through a circuit is given by i = 56.4 sin (754t)
A, where t is in seconds. What is the rms current, and what
is the frequency of the ac source?
1
Expert's answer
2020-12-21T11:02:10-0500

"I_{rms}=0.707I_m=0.707\cdot56.4\approx40(A)"


"754=2\pi\nu\to \nu=\frac{754}{2\cdot3.14}=120(Hz)"